OpenStudy (anonymous):

I have a pretty simple algebra 1 question i need answered from you guys it would be awesome if you could help Fan and Medal

OpenStudy (anonymous):

What is the side length b in the triangle below? 

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

I have 4 more triangle questions but i think ill put them up in different questions so that whoever helps me can get the right amount of medals

OpenStudy (anonymous):

i have to go ill be back in 2 minutes to look at reply

OpenStudy (anonymous):

nvm im staying i thought i had to get off

OpenStudy (mrnood):

The drawing is corrupted - the text is in the wrong place. However - I am sure you know the Pythagoras formala for the sides of a right angled triangle If oyu have 2 sides then the other is VERY simply derived from that formula

OpenStudy (anonymous):

The text is in the correct place it looks exactly like the triangle in the work there is a number on the vertical side and a number on the hypotenuse then variable is the horizontal base

OpenStudy (anonymous):

i can make it again if you need it

OpenStudy (mrnood):

in my view the text is corrupt is the vertical side 5, and the hypotenuse 13?

OpenStudy (anonymous):

Yes!! haha

OpenStudy (anonymous):

i tried to put the text as close as possible to the triangle

OpenStudy (mrnood):

so what is the Pytagoras formula? (you know I will not give you the answer - this is a basic triangle question)

OpenStudy (anonymous):

a^2 + b^2 = c^2 i think

OpenStudy (mrnood):

so - you know a and c solve for b

OpenStudy (anonymous):

13.9 right its one of my answer choices too

OpenStudy (mrnood):

that looks like a guess - and may or may not be right solve the equation

OpenStudy (anonymous):

no i did it it was 13^2 + 5^2 = sqrt 192 which is 13.9

OpenStudy (anonymous):

dont quote me on the 192 i dont exactly remember the number that was on the calculator

OpenStudy (mrnood):

it is not correct you know a and c you need to re-arrange the equation to solve for b

OpenStudy (anonymous):

How would i rearange it

OpenStudy (mrnood):

write the equation a^2+b^2 = c^2 but substitute the values you know

OpenStudy (mrnood):

write the equation as b^2 = .....

OpenStudy (anonymous):

but then its just the same thing

OpenStudy (anonymous):

12, no? pythagoras theorem

OpenStudy (anonymous):

my answer choices are 12, 144, 194, and 13.9

OpenStudy (anonymous):

@newtoos

OpenStudy (mrnood):

please try to do the things I suggest write the equation as b^2 = .....

OpenStudy (anonymous):

if i put it the way mr. nood said it would be the same because i am squaring both numbers then finding the square root

OpenStudy (mrnood):

Trust me - I will not suggest something which does not help---

OpenStudy (anonymous):

It still came out to 13.9

OpenStudy (anonymous):

fml

OpenStudy (mrnood):

You haven't done what I asked the equation is a^2 + b^2 = c^2 rewrite that so it is b^2 = ..........

OpenStudy (anonymous):

I did so that means c^2 plus a^2 is again 193. The only number that squared is 193 is 13.9

OpenStudy (mrnood):

Look - I am TRYING to help the answer is NOT 13.9 and oyu are NOT doing what iasked If oyu want th eanswer then ask someone else Ohterwise CAREFULLY do what I asked b^2 = ???

OpenStudy (anonymous):

B^2 = A^2 + c^2

OpenStudy (mrnood):

That is not correct be careful with the signs

OpenStudy (anonymous):

I need to get this question done and i have no idea what you mean when your saying it! What signs i used the only signs that are in the Pythagorean theorem

OpenStudy (mrnood):

the signs ARE in the equation, but when you changed the equation you made a mistake with the sign a^2 + b^2 = c^2 (hint: subtract a^2 from both sides of the equation) (btw - you asked me directly for help - I cannot help if you think the problem is with me)